An architect was hired to develop a modern house design. he measured the area of the house to be 700m².What is the equivalent area of the house in cm²?


1
Expert's answer
2021-09-14T09:35:23-0400
"700\ m^2\cdot(\dfrac{100\ cm}{1\ m})^2=7\cdot10^6\ cm^2."